-2x+5y=6
6x-2y=34

Answer by jim_thompson5910(35256)

Hint:

-2x+5y=6
6x-2y=34


3(-2x+5y)=3*6
6x-2y=34


-6x+15y=18
6x-2y=34
-------------
0x+13y=52

13y = 52

y = 52/13

y = 4

Now that you know that y = 4, you can use this to find x.